# DemMamba: Alignment-free Raw Video Demoiréing with Frequency-assisted Spatio-Temporal Mamba

### v2

@@ Abstract
Moiré patterns, resulting from the interference of two similar repetitive patterns, are frequently observed during the capture of images or videos on screens. These patterns vary in color, shape, and location across video frames, posing challenges in extracting information from adjacent frames and preserving temporal consistency throughout the restoration process. Existing deep learning methods often depend on well-designed alignment modules, such as optical flow estimation, deformable convolution, and cross-frame self-attention layers, incurring high computational costs. Recent studies indicate that utilizing raw data as input can significantly improve the effectiveness of video demoiréing by providing the pristine degradation information and more detailed content. However, previous works fail to design both efficient and effective raw video demoiréing methods that can maintain temporal consistency and prevent degradation of color and spatial details. This paper introduces a novel alignment-free raw video demoiréing network with frequency-assisted spatio-temporal Mamba (DemMamba). It features sequentially arranged Spatial Mamba Blocks (SMB) and Temporal Mamba Blocks (TMB) to effectively model the inter- and intra-relationships in raw videos affected by moiré patterns. An Adaptive Frequency Block (AFB) within the SMB facilitates demoiréing in the frequency domain, while a Channel Attention Block (CAB) in the TMB enhances the temporal information interactions by leveraging inter-channel relationships among features. Extensive experiments demonstrate that our proposed DemMamba surpasses state-of-the-art methods by 1.3 dB in PSNR, and also provides a satisfactory visual experience.
1 Introduction
@@
Figure 1: Motivation of our alignment-free raw video demoiréing network. (a) Most of the previous methods utilize a well-designed module for alignment. (b) We propose Spatial Mamba and Temporal Mamba for effective inter- and intra-relationship modeling.
Moiré patterns appear as repetitive, wave-like, colorful distortions when capturing images or videos on screens. These patterns result from spatial frequency aliasing between the color filter array of camera sensors and the LCD screen subpixels [39], leading to an unsatisfactory visual experience. Video demoiréing, compared to image demoiréing, poses greater challenges; it demands not only the restoration of high-quality, moiré-free frames with natural colors but also significant effort to maintain temporal consistency.
In recent years, many learning-based video demoiréing approaches [6, 34, 47, 26, 32, 5] have been proposed. Specifically, VDmoiré [6] establishes the first video demoiréing dataset captured with hand-held cameras and designs a baseline video demoiréing model that utilizes implicit feature space alignment and selective feature aggregation. DTNet [47] introduces a direction-aware, temporally guided network with bilateral learning for video demoiréing, employing both multi-scale architecture and deformable convolution for the alignment process. Many studies suggest that eliminating moiré patterns in the raw domain is more effective than in the sRGB domain [51, 5, 48], as moiré patterns in the raw domain are less pronounced due to the absence of nonlinear operations, such as gamma correction and demosaicing, typical in image signal processing (ISP), which often amplify these artifacts. Given the increasing availability of raw data in modern smartphones and DSLR cameras, demoiréing in the raw domain is both feasible and beneficial. RawVD [5] is a pioneering work that performs video demoiréing in the raw domain using a dual-branch convolution approach, which combines cross-channel and color-group convolutions to remove moiré patterns and enhance visual structures.
@@
Most of the aforementioned video demoiréing methods rely heavily on well-designed modules, such as the pyramid cascading deformable (PCD) module proposed by EDVR [42], which uses deformable convolutions to align multiple frames, as depicted in Fig. 1(a). Although this approach appears to be straightforward, it incurs a high computational burden. This poses significant challenges to efficiency, particularly in processing high-resolution, long video sequences, hampering its application in modeling long-term inter- and intra-frame correlations. Recently, State Space Models (SSMs) [18], originating from control systems theory, have demonstrated advantages in natural language processing [10, 9] and computer vision [58, 35] due to their linear complexity in handling long sequences. In low-level vision, some works have extended the application of Mamba to image restoration [12, 21, 7, 1], effectively balancing the global receptive field and computational efficiency. This prompts us to further delve into the potential of Mamba in video restoration, particularly applying it to raw video demoiréing, aiming to efficiently remove moiré patterns while maintaining temporal consistency. To facilitate effective inter- and intra-relationship modeling in raw videos with moiré patterns, we propose the Spatial Mamba and Temporal Mamba modules, which are arranged sequentially, as demonstrated in Fig. 1(b).
@@
In this paper, we propose an alignment-free raw video demoiréing network, termed DemMamba, featuring frequency-assisted spatio-temporal Mamba. Our principal design employs sequentially arranged Spatial Mamba Blocks (SMB) and Temporal Mamba Blocks (TMB) to effectively model inter- and intra-relationships, thereby efficiently removing moiré patterns while maintaining temporal consistency. For SMB and TMB, we adopt different scanning strategies and introduce additional modules to enhance correlations within the spatial and temporal domains, respectively. Considering the formation process of moiré patterns, we utilize a learnable compressor in the frequency domain to selectively attenuate frequencies associated with prominent moiré patterns in SMB. For TMB, channel attention is employed to further enhance temporal information interactions by exploiting inter-channel relationships among features.
@@
    We adapt Mamba into a spatio-temporal framework for the video restoration task and design a frequency-assisted spatio-temporal Mamba network for raw video demoiréing. By alternately employing Spatial Mamba and Temporal Mamba blocks, DemMamba effectively alleviates moiré patterns and maintains temporal consistency in an alignment-free approach.
@@
    SMB is utilized for the intra-frame demoiréing process. We introduce a learnable compressor to attenuate prominent moiré frequencies in the frequency domain.
@@
    TMB is designed to capture inter-channel relationships, which facilitates better utilization of neighboring information and maintains temporal consistency. The temporal sequences are processed through simultaneous forward and backward SSMs, and channel attention is introduced to enhance the integration of temporal information.
  • $\bullet$
@@
    We conduct extensive experiments on both raw video and image demoiréing datasets. DemMamba surpasses state-of-the-art methods by 1.3 dB in PSNR, and also delivers a visually appealing result.
2 Related Works
2.1 Image and Video Demoiréing
@@
Moiré patterns result from the interference between two similar frequency patterns, frequently appearing in screen captures and severely degrading the visual experience. Several learning-based solutions have been proposed to eliminate moiré patterns from images [37, 23, 13, 4, 28, 25, 24, 14, 57, 40, 50, 31, 39, 55, 48, 45]. Compared to image demoiréing, video demoiréing [6, 34, 47, 32] presents a greater challenge, as it requires generating temporally consistent moiré-free predictions. VDMoiré [6] introduces a straightforward video demoiréing model that leverages implicit feature space alignment and selective feature aggregation. Additionally, it incorporates a novel relation-based consistency loss to enhance temporal consistency across frames. DTNet [47] proposes a direction-aware, temporally guided bilateral learning network for video demoiréing, utilizing multi-scale alignment and deformable convolution during the alignment process. RawVD [5] employs a pyramid cascading deformable module for the alignment process and utilizes a dual-branch convolution approach that combines cross-channel and color-group convolutions to effectively remove moiré patterns and enhance visual structures. Notably, existing video demoiréing methods heavily rely on sophisticated alignment modules, which significantly increasing computational complexity. Therefore, we introduce DemMamba, an alignment-free method designed to effectively and efficiently eliminates moiré patterns from videos, maintaining temporal consistency while reducing computational demands.
2.2 Deep Raw Image and Video Restoration
@@
Raw pixels provide richer information for visual tasks due to their broader bit depth, as they inherently contain a vast amount of data. Recently, several low-level vision tasks have adopted raw data, including low-light enhancement [15, 8, 16], super-resolution [46, 52, 30], reflection removal [20, 36], and moiré pattern removal [51, 48, 5]. RRID [48] proposes an image demoiréing network that employs paired raw-sRGB data to facilitate the color recovery process. RawVD [5], the first work to explore raw video demoiréing, constructs the RawVDemoiré dataset. In this paper, considering that raw pixels offer more information and that moiré patterns are less apparent in the raw domain, we perform video demoiréing in the raw domain.
2.3 State Space Models
@@
State Space Models (SSMs) [18] have garnered significant attention recently due to their notable efficiency in employing state space transformations to manage long-term dependencies in language sequences [11]. More recently, Mamba [9] is designed with a selective scan mechanism and an efficient hardware architecture, demonstrating substantial performance improvements. These advancements validate the effectiveness of Mamba in visual tasks, extending its applications to semantic segmentation [27, 59, 38], image classification [29, 58, 33, 49, 2], video understanding [41, 3], and image restoration [12, 56, 44, 21, 22, 19]. Building on Mamba’s success in these areas, this paper explores the potential of Mamba in raw video demoiréing. The frequency-assisted spatio-temporal design of our proposed DemMamba also presents an effective approach that could be adapted for other video restoration tasks.
@@
3 Proposed Method
Moiré patterns, initially caused by frequency interference, are further exacerbated by the nonlinear processes in the ISP. Thus, introducing raw data for video demoiréing may provide greater effectiveness. The color, shape, and location of moiré patterns can vary across different frames due to changing camera viewpoints, automatic focal length adjustments, and other factors. Therefore, the demoiréing model must capture long-range dependencies in both spatial and temporal domains. This capability is essential for effectively leveraging global information from the current frame and relevant temporal information from adjacent frames to remove moiré patterns while ensuring temporal consistency across the video sequence.
In this work, we propose a frequency-assisted spatio-temporal Mamba for raw video demoiréing. The Spatial Mamba Block (SMB) and Temporal Mamba Block (TMB) are sequentially arranged to model intra- and inter-relationships in raw videos affected by moiré patterns.
3.1 Preliminaries
The recent advancements in structured state-space sequence models (S4) are largely inspired by the continuous linear time-invariant (LTI) systems, which map a 1-dimensional input function or sequence $x(t)\in\mathbb{R}$ to an output $y(t)\in\mathbb{R}$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$. Formally, this system can be described by a linear ordinary differential equation (ODE):
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t)+\mathbf{D}x(t),$
where $N$ denotes the state size, $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times N}$, and $\mathbf{D}\in\mathbb{R}$.
Subsequently, the discretization process is typically employed to integrate Eq. 1 within practical deep learning algorithms. Specifically, let $\Delta$ denote the timescale parameter to transform the continuous parameters $\mathbf{A}$, $\mathbf{B}$ to discrete parameters $\overline{\mathbf{A}}$, $\overline{\mathbf{B}}$. The most commonly used method for discretization is the zero-order hold (ZOH) rule, which is formulated as:
$\overline{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),$ (2)
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.$
After the discretization, the discrete version of Eq. 1 with step size $\Delta$ can be rewritten in the following RNN form:
$h_{k}$ $=\overline{\mathbf{A}}h_{k-1}+\overline{\mathbf{B}}x_{k},$ (3)
$y_{k}$ $=\mathbf{C}h_{k}+\mathbf{D}x_{k}.$
Furthermore, Eq. 3 can also be equivalently transformed into the following CNN form mathematically:
$\overline{\mathbf{K}}$ $\triangleq(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\cdots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),$ (4)
$\mathbf{y}$ $=\mathbf{x}\circledast\overline{\mathbf{K}},$
[…]
Building on S4’s framework, the recent advanced state-space model, Mamba [9], has further improved $\overline{\mathbf{B}}$, $\mathbf{C}$, and $\Delta$ to be input-dependent, thereby enabling dynamic feature representations. Mamba’s approach to image restoration leverages the same recursive form of Eq. 3, allowing the model to capture long sequences, thus aiding in pixel-wise restoration. Additionally, the parallel scan algorithm [9] allows Mamba to enjoy the same advantages of parallel processing as Eq. 4, thus facilitating efficient training.
3.2 Architecture
@@
For the preprocessing, the raw moiré images $\mathbf{I}^{t+i}_{raw}\in\mathbb{R}^{H\times W\times 3}$, $i\in\{{-}1,0,{+}1\}$, where $H\times W$ represents the spatial resolution, are initially packed into the 4-channel RGGB format $\mathbf{I}^{t+i}_{in}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$, $i\in\{{-}1,0,{+}1\}$. To address various moiré patterns across frames, we concatenate the packed inputs $\mathbf{I}^{t+i}_{in},i\in\{{-}1,0,{+}1\}$, and perform shallow feature extraction to obtain the shallow features $\mathbf{F}_{0}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$ through a single convolutional layer with a stride of two. Subsequently, we employ several Frequency-assisted Spatio-Temporal Mamba Groups (FASTMG) to learn the spatio-temporal features. Within each FASTMG, spatial and temporal dependencies are handled by specific modules. Each FASTMG consists of $M$ SMB and $M$ TMB organized sequentially to model both inter- and intra-frame relationships effectively, followed by a convolutional layer. In SMB, a 2-D selective scan is employed to capture spatial features, providing both a global effective receptive field and linear computational complexity relative to the input size. In conjunction with the spatial Mamba, we introduce an Adaptive Frequency Module (AFM) with an adaptive compressor in the frequency domain to attenuate the specific frequencies of moiré patterns. TMB processes flattened temporal sequences through simultaneous forward and backward SSMs, enhancing its capacity for temporal-aware processing. A Channel Attention Block (CAB) is embedded to further enhance temporal information interactions by exploiting inter-channel relationships among features. To preserve and integrate shallow and spatio-temporal features, we incorporate a global residual connection. Finally, the reconstruction module derives $\mathbf{\hat{I}}^{t}_{c}$. The reconstruction module consists of several convolutional layers and a sub-pixel convolution layer to upsample the features, thus reconstructing the predicted clean image with height $H$ and width $W$.
@@
3.2.1 Design of SMB
SMB, depicted in Fig. 2(a), is designed to establish spatial relationships of moiré patterns. Inspired by the success of Mamba in long-range modeling with linear complexity, we introduce the Spatial Mamba, shown in Fig. 2(c), to capture long-range dependencies in the spatial domain.
Following the standard Vision SSM design [29], the input feature $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is processed through two parallel branches. In the first branch, the feature channels are expanded to $\lambda C$ by a linear layer, where $\lambda$ is a pre-defined channel expansion factor, followed by depth-wise convolution, the SiLU activation function, the 2D-SSM layer, and LayerNorm. In the second branch, the feature channels are similarly expanded to $\lambda C$ using a linear layer, then processed with the SiLU activation function. Afterward, features from both branches are aggregated using the Hadamard product. The final output of the Spatial Mamba $\mathbf{X}_{SM}$ is obtained by projecting the channel number back to $C$. The process in Spatial Mamba can be described as:
@@ 3.2.1 Design of SMB
To extend selective SSM to 2D spatial data, we follow [29] and introduce a 2D Selective Scan Module for cross scanning in Spatial SSM. As illustrated in Fig. 2(c.1), the 2D image feature is flattened and scanned along four directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Long-range dependencies are captured in each sequence using discrete state-space equations, and the sequences are then summed and reshaped to restore the 2D structure.
Considering that moiré patterns arise from frequency aliasing, we propose AFB to aid in the removal of moiré patterns in the frequency domain, operating in parallel with the Spatial Mamba. It is assumed that the frequency spectrum of moiré patterns tends to be relatively consistent within small patches. In signal processing, a band-reject filter allows most frequencies to pass unaltered while attenuating those within a specific range to very low levels. Given that different patches may require the attenuation of distinct frequencies, identifying a frequency prior for each moiré image patch is time-consuming. Consequently, as shown in Fig. 3(a), we construct a learnable instance-adaptive compressor $C(\cdot)$ in AFM, composed of several convolutional layers, to dynamically target and filter frequencies associated with moiré patterns while preserving the original image content. We use Real FFT and its inverse process to apply $C(\cdot)$ in the frequency domain as:
@@
$\hat{\mathbf{X}}=F^{-1}((C(\mathbf{X}_{F}))\odot\mathbf{X}_{F}),$
where $F(\cdot)$ and $F^{-1}(\cdot)$ represent Real FFT and its inverse process, respectively.
Finally, the output of the SMB is produced through a weighted summation of the outputs from the Spatial Mamba and the AFB as:
$\mathbf{X}_{SMB}=\mathbf{X}_{SM}+\alpha_{1}\mathbf{X}_{F},$ (7)
where $\alpha_{1}$ is a pre-defined weighting coefficient that enables the model to effectively reduce moiré patterns while maintaining overall visual quality and detail richness.
Table 1: Comparison with state-of-the-art image and video demoiréing methods for raw video demoiréing in terms of average PSNR, SSIM, LPIPS, and computing complexity. The best result is highlighted in bold and the second best is underlined.

### v3

# Alignment-free Raw Video Demoiréing

@@ Abstract
Video demoiréing aims to remove undesirable interference patterns that arise during the capture of screen content, restoring artifact-free frames while maintaining temporal consistency. Existing video demoiréing methods typically utilize carefully designed alignment modules to estimate inter-frame motion for leveraging temporal information; however, these modules are often complex and computationally demanding. Meanwhile, recent works indicate that using raw data as input significantly enhances demoiréing performance. Building on this insight, this paper introduces a novel alignment-free raw video demoiréing network with frequency-assisted spatio-temporal Mamba (DemMamba). It incorporates sequentially arranged Spatial Mamba Blocks (SMB) and Temporal Mamba Blocks (TMB) to effectively model the inter- and intra-relationships in raw video demoiréing. The SMB employs a multi-directional scanning mechanism coupled with a learnable frequency compressor to effectively differentiate interference patterns across various orientations and frequencies, resulting in reduced artifacts, sharper edges, and faithful texture reconstruction. Concurrently, the TMB enhances temporal consistency by performing bidirectional scanning across the temporal sequences and integrating channel attention techniques, facilitating improved temporal information fusion. Extensive experiments demonstrate that DemMamba surpasses state-of-the-art methods by 1.6 dB in PSNR, and also delivers a satisfactory visual experience.
Introduction
@@
Figure 1: Motivation of our alignment-free raw video demoiréing network. (a) Previous methods utilize a well-designed alignment module. (Computational unfriendly.) (b) We propose Spatial Mamba and Temporal Mamba for inter- and intra-relationship modeling. (Effective and efficient.)
Moiré patterns appear as repetitive, wave-like, and colorful distortionsin images or videos captured from screens. These patterns result from frequency aliasing between the color filter array of camera sensors and the LCD screen subpixels (Wang et al. 2023a), resulting in a degraded visual experience. Image demoiréing has received increasing attention from the research community (Cheng, Fu, and Yang 2019; He et al. 2020; Zheng et al. 2020; Wang et al. 2021; Niu et al. 2023; Wang et al. 2023a). In these image demoiréing methods, the restoration process can be modeled as a combination of moire pattern removal and color correction. Many works indicate that addressing moiré patterns early in the raw-data pipeline yields superior results compared to corrections performed post-Image Signal Processor (ISP) in the sRGB domain (Yue et al. 2022; Cheng, Liu, and Yang 2024; Xu et al. 2024a), as the raw domain is free of nonlinear processes (e.g., gamma correction and demosaicing) that exacerbate moiré artifacts. Given the increasing availability of raw data in smartphones and DSLR cameras, demoiréing in the raw domain is both practical and beneficial.
With the growing demand for video capture, several learning-based methods have been developed to mitigate undesirable moiré patterns in videos (Dai et al. 2022; Quan et al. 2023; Xu et al. 2024b; Liu et al. 2024a; Niu et al. 2024; Cheng, Liu, and Yang 2024). However, video demoiréing presents greater complexities compared to image demoiréing. It demands not only the restoration of high-quality, moiré-free frames with natural colors but also significant effort to maintain temporal consistency.
To leverage information from adjacent frames and maintain temporal consistency, most of the existing video demoiréing methods rely heavily on sophisticated alignment modules, such as the pyramid cascading deformable (PCD) module proposed by EDVR (Wang et al. 2019). This module employs deformable convolutions to align multiple frames, as depicted in Fig. 1(a). Although this approach appears to be straightforward, it incurs a high computational burden, limiting its efficiency, especially for processing high-resolution, long-duration videos.
@@
Recently, State Space Models (SSMs) (Kalman 1960), originating from control system theory, have demonstrated advantages in natural language processing(Gu, Goel, and Ré 2021; Gu and Dao 2023) and computer vision (Zhu et al. 2024a; Shi et al. 2024), primarily due to their linear complexity in processing long sequences. In low-level vision tasks, some works have successfully adapted the Mamba architecture for image restoration (Guo et al. 2025; Li et al. 2024; Deng and Gu 2024; Bai et al. 2024), effectively balancing global receptive fields and computational efficiency. This prompts us to further delve into the potential of Mamba in video restoration, particularly targeting efficient removal of moiré patterns from raw videos while preserving temporal consistency. To facilitate modeling effective inter-frame and intra-frame relationship in raw videos with moiré patterns, as demonstrated in Fig. 1(b), we propose the sequentially arranged Spatial Mamba and Temporal Mamba blocks.
@@
In this paper, we propose an alignment-free raw video demoiréing network, DemMamba, featuring frequency-assisted spatio-temporal Mamba. Our principal design employs sequentially arranged Spatial Mamba Blocks (SMB) and Temporal Mamba Blocks (TMB) to model inter- and intra-relationships, thereby efficiently removing moiré patterns while maintaining temporal consistency. Considering that the orientation of moiré patterns can vary substantially even within a single image, we introduce a multi-directional scanning mechanism within the SMB. This mechanism minimizes directional bias, accommodates diverse textural interferences, and enhances moiré artifact suppression. Additionally, we integrate a learnable frequency-domain compressor within the SMB to selectively attenuate spectral components associated with pronounced moiré patterns. For TMB, the bidirectional scanning effectively models inter-frame relationships, enhancing the utilization of neighboring information. Furthermore, we integrate a channel attention block within the TMB to reinforce inter-channel dependencies, thereby further suppressing moiré patterns and ensuring robust temporal consistency throughout the video sequence.
@@
    We propose an alignment-free framework for raw video demoiréing. It alternately applies Spatial Mamba and Temporal Mamba blocks to model intra- and inter-frame relationships, effectively removing moiré patterns while preserving temporal consistency.
@@
    For raw video demoiréing, we propose the SMB embedded with a multi-direction scanning mechanism and a learnable frequency-domain compressor to suppress prominent moiré patterns. Also, we introduce the TMB to model inter-frame relationships, exploiting neighboring information and ensuring temporal consistency.
@@
    Extensive experiments are conducted on both raw video and image demoiréing datasets. DemMamba surpasses state-of-the-art methods by 1.6 dB in PSNR, and also delivers a visually appealing result.
Related Works
Image and Video Demoiréing
@@
Moiré patterns result from the interference between two similar frequency patterns, frequently appearing in screen captures and severely degrading the visual experience. Several learning-based solutions have been proposed to eliminate moiré patterns from images (Liu, Shu, and Wu 2018; He et al. 2019; Liu et al. 2020c, b; Zhang et al. 2023; Xiao, Lu, and Wang 2020; Yang et al. 2025). Compared to image demoiréing, video demoiréing (Dai et al. 2022; Quan et al. 2023; Xu et al. 2024b; Niu et al. 2024) presents a greater challenge, as it requires generating temporally consistent moiré-free predictions. VDMoiré (Dai et al. 2022) introduces a straightforward video demoiréing model that leverages implicit feature space alignment and selective feature aggregation. Additionally, it incorporates a novel relation-based consistency loss to enhance temporal consistency across frames. DTNet (Xu et al. 2024b) proposes a direction-aware, temporally guided bilateral learning network for video demoiréing, utilizing multi-scale alignment and deformable convolution during the alignment process. RawVD (Cheng, Liu, and Yang 2024) employs a pyramid cascading deformable module for the alignment process and utilizes a dual-branch convolution approach that combines cross-channel and color-group convolutions to effectively remove moiré patterns and enhance visual structures. Notably, existing video demoiréing methods rely on complex alignment modules, increasing computational complexity. Therefore, we introduce DemMamba, an alignment-free method designed to effectively and efficiently eliminates moiré patterns from videos, maintaining temporal consistency while reducing computational demands.
Deep Raw Image and Video Restoration
@@
Raw pixel data inherently offer richer information for visual tasks, owing to their greater bit depth and extended radiance range. Recently, several low-level vision tasks have adopted raw data, including low-light enhancement (Huang et al. 2022; Dong et al. 2022; Jin et al. 2023), super-resolution (Xing and Egiazarian 2021; Yue, Zhang, and Yang 2022; Luo, Wu, and Guo 2024), reflection removal (Lei and Chen 2021; Song et al. 2023), and moiré pattern removal (Yue et al. 2022; Xu et al. 2024a; Cheng, Liu, and Yang 2024). RRID (Xu et al. 2024a) proposes an image demoiréing network that employs paired raw-sRGB data to facilitate the color recovery process. RawVD (Cheng, Liu, and Yang 2024), the first work to explore raw video demoiréing, constructs the RawVDemoiré dataset. In this paper, considering that raw pixels offer more information and that moiré patterns are less apparent in the raw domain, we perform video demoiréing in the raw domain.
State Space Models
@@
State Space Models (SSMs) (Kalman 1960) have garnered significant attention recently due to their notable efficiency in employing state space transformations to manage long-term dependencies in language sequences (Gu et al. 2021). More recently, Mamba (Gu and Dao 2023) is designed with a selective scan mechanism and an efficient hardware architecture, demonstrating substantial performance improvements. These advancements validate the effectiveness of Mamba in visual tasks, extending its applications to semantic segmentation (Liu et al. 2024b; Zhu et al. 2024b; Wan et al. 2024), image classification (Liu et al. 2024c; Zhu et al. 2024a; Patro and Agneeswaran 2024; Yang et al. 2024; Behrouz, Santacatterina, and Zabih 2024), video understanding (Wang et al. 2023b; Chen et al. 2024), and image restoration (Guo et al. 2025; Zhen, Hu, and Feng 2024; Wu et al. 2024; Li et al. 2024; Lin et al. 2024; Kong et al. 2024). Building on Mamba’s success, this paper explores the potential of Mamba in raw video demoiréing.
Proposed Method
Motivation
Utilizing raw data. We advocate using raw data for video demoiréing, as moiré patterns are more distinguishable from content in the raw domain because the raw data remains unaffected by nonlinear operations in the ISP (Cheng, Liu, and Yang 2024). This distinction is clearly illustrated in Fig.2, where the Discrete Cosine Transform (DCT) is applied to images containing moiré patterns in both raw and sRGB domains. The raw domain exhibits distinct frequency distributions across color channels, whereas the DCT spectra for the R, G, and B channels in the sRGB domain show higher similarity, making moiré patterns less discernible. Therefore, we prefer raw data over sRGB data for video demoiréing.
Figure 2: DCT analysis of the moiré raw and sRGB images.
Applying Mamba-based framework. Mamba (Gu and Dao 2023) incorporates a selective scan mechanism and exhibits linear complexity when processing long sequences, making it intuitively suitable for video restoration tasks. Building upon the foundation established by Mamba, we introduce a novel configuration of Spatial Mamba Blocks (SMBs) and Temporal Mamba Blocks (TMBs), designed to effectively model both inter-frame and intra-frame relationships for video demoiréing. SMB is designed to produce high-quality, moiré-free video frames with natural color fidelity. Meanwhile, TMB leverage temporal data from adjacent frames to remove moiré patterns while preserving temporal consistency across the video sequence.
@@
Architecture
@@
For the preprocessing, the raw moiré images $\mathbf{I}^{t+i}_{raw}\in\mathbb{R}^{H\times W\times 3}$, $i\in\{{-}1,0,{+}1\}$, where $H\times W$ represents the spatial resolution, are initially packed into the 4-channel RGGB format $\mathbf{I}^{t+i}_{in}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$. To address various moiré patterns across frames, we concatenate the packed inputs $\mathbf{I}^{t+i}_{in}$, and perform shallow feature extraction to obtain the shallow features $\mathbf{F}_{0}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$ through a single convolutional layer with a stride of two , where $C$ is the number of feature channels. Subsequently, we employ several Frequency-assisted Spatio-Temporal Mamba Groups (FASTMG) to learn the spatio-temporal features. Within each FASTMG, spatial and temporal dependencies are handled by specific modules. Each FASTMG consists of $M$ SMBs and $M$ TMBs organized sequentially to model both inter- and intra-frame relationships effectively, followed by a convolutional layer. In SMB, a multi-direction scan is employed to capture spatial features in the Spatial Mamba, providing both a global effective receptive field and linear computational complexity relative to the input size. In conjunction with the Spatial Mamba, we introduce an Adaptive Frequency Block (AFB) with an adaptive compressor in the frequency domain to attenuate the specific frequencies of moiré patterns. TMB processes flattened temporal sequences through simultaneous forward and backward SSMs, enhancing its capacity for temporal-aware processing. A Channel Attention Block (CAB) is embedded to further enhance temporal information interactions by exploiting inter-channel relationships among features. To preserve and integrate shallow and spatio-temporal features, we incorporate a global residual connection. Finally, the reconstruction module derives $\mathbf{\hat{I}}^{t}_{c}$. The reconstruction module consists of several convolutional layers and a sub-pixel convolution layer to upsample the features, thus reconstructing the predicted clean image with height $H$ and width $W$.
@@
Design of SMB
SMB, depicted in Fig. 3(a), is designed to establish spatial relationships among moiré patterns, facilitating the production of high-quality, moiré-free video frames with authentic color fidelity. Inspired by the success of Mamba in long-range modeling with linear complexity, we introduce the Spatial Mamba, shown in Fig. 3(c), to capture long-range dependencies in the spatial domain.
Following the Vision SSM design (Liu et al. 2024c), the input feature $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is processed through two parallel branches. In the first branch, the feature channels are expanded to $\lambda C$ by a linear layer, where $\lambda$ is a pre-defined channel expansion factor, followed by depth-wise convolution, the SiLU activation function, the 2D-SSM layer, and LayerNorm. In the second branch, the feature channels are similarly expanded to $\lambda C$ using a linear layer, then processed with the SiLU activation function. Afterward, features from both branches are aggregated using the Hadamard product. The final output of the Spatial Mamba $\mathbf{X}_{SM}$ is obtained by projecting the channel number back to $C$. The process in Spatial Mamba can be described as:
@@
To extend selective SSM to 2D spatial data, we introduce a 2D Selective Scan Module for multi-direction scanning in Spatial SSM. Motivated by the fact that moire patterns often exhibit diverse directional characteristics, conventional single-directional scanning methods are insufficient for accurately modeling and removing these complex artifacts. By scanning along multiple orientations, the module can comprehensively capture moiré patterns that appear at arbitrary angles, ensuring robust detection of interference edges regardless of their direction. As shown in Fig. 3(c.1), the 2D image feature is flattened and scanned along eight directions. Long-range dependencies are captured in each sequence using discrete state-space equations, and the sequences are summed and reshaped to restore the 2D structure.
Considering that moiré patterns arise from frequency aliasing, we propose AFB to aid in the removal of moiré patterns in the frequency domain, operating in parallel with the Spatial Mamba. It is assumed that the frequency spectrum of moiré patterns tends to be relatively consistent within small patches (Xu et al. 2024b). In signal processing, a band-reject filter allows most frequencies to pass unaltered while attenuating those within a specific range to very low levels. Given that different patches may require the attenuation of distinct frequencies, identifying a frequency prior for each moiré image patch is time-consuming. Consequently, as shown in Fig. 4(a), we construct a learnable instance-adaptive compressor $L(\cdot)$ in AFB, composed of several convolutional layers, to dynamically target and filter frequencies associated with moiré patterns while preserving the original image content. We use Real FFT and its inverse process to apply $L(\cdot)$ in the frequency domain as:
@@
$\hat{\mathbf{X}}=F^{-1}((L(\mathbf{X}_{F}))\odot\mathbf{X}_{F}),$
where $F(\cdot)$ and $F^{-1}(\cdot)$ represent the Real FFT and its inverse process, respectively.
Finally, the output of the SMB is produced through a weighted summation of the outputs from the Spatial Mamba and the AFB as $\mathbf{X}_{SMB}=\mathbf{X}_{SM}+\alpha_{1}\mathbf{X}_{F}$, where $\alpha_{1}$ is a pre-defined weighting coefficient that enables the model to effectively reduce moiré patterns while maintaining overall visual quality and detail richness.